• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jroche, cshloka

    Log entry time 00:13:36 on October 06, 2010

    Entry number 333765

    This entry is a followup to: 333685

    keyword=BPMs have been calibrated

    The BPMs have been calibrated based on the data taken yesterday (entry 333685) and the night 
    before (entry 333567).   The runs  taken during those nights have been re-analyzed with the 
    new calibration to check that the calibration works, it does. I also did update the spot++ 
    database (/adaqfs/home/adaq/rastersize/LeftHRS/DB), I hope that was the good thing to do.
    
    As instructed we calibrated the bpms coda readings against the coda beam reading. This said 
    we noticed that the harps readings and the coda readings in the Y directions have a ~1.5 mm 
    offset.
    
    For example
    6964: epics,harps
    AX:-0.05,-0.14  AY:0.02,1.75
    BX:0.00,0.10      BY:0.00,1.45
    6969:epics-harps
    AX:-3.0,-3.1 AY:-3.1,-1.6
    BX:-3.1,-3.1 BY:-2.9,-1.6
    
    
    
    


    A copy of this log entry has been emailed to: cshloka@jlab.org